OpenStudy (theqube1):

I need help with this Geometry PLEASE. I'm not asking for just the answer but how to go about finding it. I will Fan and Medal Best responders

OpenStudy (here_to_help15):

Ok

OpenStudy (anonymous):

What's the question?

OpenStudy (theqube1):

Here is the question @lexiladybug22 @Here_to_Help15

OpenStudy (here_to_help15):

Connections?

OpenStudy (theqube1):

Yes

OpenStudy (theqube1):

@Here_to_Help15 are you in geometry B? and do you know this lesson?

OpenStudy (here_to_help15):

Ok good um i need help actually ;) and yes i am

OpenStudy (theqube1):

Dang lol you on the final to?

OpenStudy (here_to_help15):

Nope

OpenStudy (theqube1):

oh well that is what this is.

ganeshie8 (ganeshie8):

which part are you working on ?

OpenStudy (theqube1):

@ganeshie8 all three

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

check whether the below proportion holds\[\large \dfrac{AC}{EC}~~\stackrel{?}{=}~~\dfrac{BC}{DC}\]

ganeshie8 (ganeshie8):

\[\large \dfrac{12}{9}~~\stackrel{?}{=}~~\dfrac{20}{15}\] \[\large \dfrac{4}{3}~~\stackrel{?}{=}~~\dfrac{4}{3}\] which is true, therefore \(\triangle ABC \sim \triangle EDC\) is true

OpenStudy (mathstudent55):

Just like there is SAS for proving two triangles congruent, there is a version of SAS for proving two triangles similar. SAS Similarity is: if the lengths of two sides of a triangle are proportional to corresponding parts of another triangle, and the included angles are congruent, then the triangles are similar.

OpenStudy (mathstudent55):

Above, @ganshie8 showed that the lengths of the sides are proportional. From the figure, you see that angles ACB and ECD are vertical, making them congruent. This is what you need for SAS Similarity to work, and the triangles are similar.

OpenStudy (theqube1):

@mathstudent55 so using SAS similarity is ow I find out that problem?

OpenStudy (mathstudent55):

Using SAS Similarity allows you to answer part (A). You can say that the triangles are similar, and the reason is SAS Similarity, since there is a pair of congruent corresponding angles, and this pair of angles are included by two pairs of corresponding sides whose lengths are proportional. All of this is part (A).

OpenStudy (theqube1):

what about part (B) and (C)? @mathstudent55